5/x+6+2/x^2+7x+6=3/x+1
This one expression is stumping me! Do I multiply through with 6?
Please help!
Thanks~~
CH

Answer by stanbon(75887)

5/(x+6) +2/(x^2+7x+6)=3/(x+1)
The LCD is x^2+7x+6
Multiply thru by the LCD to get:
5(x+1) + 2 = 2(x+6)
5x+ 5 + 2 = 2x+12
3x=5
x=5/3
Cheers,
Stan H.